OpenStudy (katienic):

how to do sig fig problem [(286.8×105)−(1.300×103)]×2.8965

OpenStudy (cuanchi):

[(286.8×105)−(1.300×103)]×2.8965 4sf 3sf 4sf 3sf 5sf Multiplication and Division The LEAST number of significant figures in any number of the problem determines the number of significant figures in the answer. 286.8×105 = 30114 = 3.01 x 10^4 4sf 3sf 3sf you can mark the last significant figure and keep one extra for the calculations 1.300 ×103 = 133.9 = 134 4sf 3sf 3sf (rounded) ----------------------------------- Addition and Subtraction For addition and subtraction, look at the decimal portion in your case both have none decimal figures 30114 - 134 =29980 no decimal places ------------------------------ 29980 × 2.8965 = 86837.07 = 86837 5sf 5sf 5sf

OpenStudy (joannablackwelder):

Are the 5 and 3 supposed to be exponents?

Nnesha (nnesha):

yes.^

Nnesha (nnesha):

\[[(286.8 \cdot 10^5) -1.300 \cdot 10^3)](2.8965)\] like this @katienic ? type `^` to represent an exponent

OpenStudy (cuanchi):

If that is the case that change everything!!!! for addition and subtraction with scientific notation the numbers has to have the same exponential term in your case both have to have 10^5 (286.8 x 10^5) - (0.01300 x 10^5) 286.8 x10^5 - 0.01300 x10^5 -------------- 268.78700 x10^5 keep the number of decimals as the lowest and rounded 268.8 x 10^5 basically it is the same number than the first in your calculation because the uncertainty of this number is bigger than the value that we want to subtract from it.